a magician pulls a tablecloth out from under some dishes. How far do the dishes move during the 0.22 it takes to pull out the tablecloth? The coefficient of kinetic friction between the cloth and the dishes is = 0.15.
Answer #1
Friction = umg = 0.15*m*9.8

F =ma
=> a = 9.8*0.15 = 1.47m/s^2

S = ut +.5at^2
u =0
t =0,22
a= 1.47
s = 0+ .5*1.47*0.22^2 =0.035574m

the dishes move 3.5574cm